The width of a rectangle is 3 meters more than one fourth its length. The perimeter is 26 meters. Find the length and width.
Answer by askme(9)

(W)idth = 3 + 1/4 (L)ength
(L)ength = L


2*(3 + (1/4)L + L) = 26 meters
1/4L + L = 26/2 - 3
1/4L + 4/4L = 10
5/4L = 10
L = 10 * 4/5
L = 40/5 = 8 meters


W = 3 + 1/4 L
W = 3 + 1/4 *8
W = 3 + 2
W = 5 meters


Length = 8 meters
Width = 5 meters